Technical Instrument Report CDBS 2008-02

Delivery of Reference Files to the Data Management Systems

R. I. Diaz-Miller, M. Cracraft rmiller@stsci.edu,cracraft@stsci.edu

January 11, 2008

Abstract

This TIR describes the INS/CDBS Team’s responsibilities. It defines the standard procedures for the delivery of calibration pipeline and SYNPHOT reference files to the Data Management Systems (DMS). It provides guidelines for test and validation of reference files by the INS/CDBS Team. This is an update to TIR CDBS 2005-02A. It clarifies procedures and describes new tools developed to facilitate the delivery process and better test the reference files. It also provides information on the new disk locations for the reference files and the new procedures adopted for the delivery of the SYNPHOT reference files.

Introduction

In order for reference files to be used by the OTFR pipeline, these files have to be copied to OPUS disks. Pointers to these files are based on instrument modes and applicability dates. On the other hand, their ingestion in the Data Archive and Distribution Services (DADS), disk media and database, allows users to retrieve them from the archive. The main function of the CDBS database is to allow selection of the correct reference files based on an instrument’s configuration and date. The selection is based on data file keyword values and criteria outlined in ICD-47 (http://www.stsci.edu/instruments/observatory/cdbs/documents/icd-47RevF.pdf or http://www.ess.stsci.edu/projects/distribution/ICD47/ICD-47 RevF.pdf.). More information can be found in the CDBS Documents web page (http://www.stsci.edu/hst/observatory/cdbs/documents).

The person creating the file will be checking and assessing the quality of the reference files. The official delivery of the files to CDBS is handled by the INS/CDBS Team. Here we describe detailed steps and procedures for ingesting reference files into the databases.